# Design and Implementation of a Scalable Financial Exchange in the Public Cloud

### v8

@@
Abstract
Financial exchanges are migrating to the cloud, but the best-effort nature of the public cloud is at odds with the stringent latency requirements of exchanges. We present Jasper, a system for meeting the networking requirements of financial exchanges on the public cloud. Jasper uses an overlay tree to scalably multicast market data from an exchange to ~1000 participants with low latency () and difference in data reception time between any two participants. Jasper reuses the same tree for scalable inbound communication (participants to exchange), augmenting it with order pacing and a new priority queue, Limit Order Queue (LOQ), to efficiently handle bursts of market orders. Jasper achieves better scalability and 50% lower latency than the AWS multicast service [1]. During bursty market activity, LOQ nearly doubles the order processing rate.
1 Introduction
@@
Financial exchanges are migrating to the public cloud for several reasons such as improved scalability and reduced capital expenditure. Despite its benefits, the cloud poses unique challenges. In particular, exchanges have traditionally operated in on-premise or colocation facilities, engineered for deterministic and low latency. For fair market access, exchanges provide deterministic latency by equalizing cable lengths between the exchange and participant servers; thus, all participants receive market data simultaneously. To reduce latency, exchanges often use FPGAs as gateways [2] and L1 switches that perform low-latency physical-layer multicast [3].
@@
However, the public cloud lacks these enhancements. It is a best-effort environment characterized by non-determinism (e.g., latency variance in Figure 1). In response, several prior projects have developed techniques for cloud-based exchanges. These include using synchronized clocks to compensate for non-determinism [4], using SmartNICs [5], and new definitions of fairness [6, 7]. These projects show promising results for 10s of participants. Our paper develops new networking techniques for scale: how do we architect an exchange to support communication between the exchange and ~1000 participants in the cloud? In practice, a scalable cloud exchange would also need scale-out compute techniques for the exchange server—a concern that is out of scope.
@@
Our system, Jasper, tackles 3 challenges to provide network support for scalable exchanges. First, how do we support outbound communication of market data from the exchange to ~1000 market participants, while ensuring (a) low spatial variance, i.e., all participants receive market data nearly simultaneously, (b) low latency from the exchange server to the participant, and (c) low temporal variance, i.e., latency doesn’t fluctuate over time? Second, how do we support inbound communication of participant orders to the exchange—especially during intense market activity when bursts of orders from many participants arrive at once, causing incast-style [8] drops and increased order latency? Third, trust: a rational participant will act on market data as soon as it is available. Hence, for fairness, the network must hold data and release it to a given participant only when it is sure that the data is available to all participants. How do we mitigate this additional holding latency and latency variance? In tackling these challenges, we make 3 contributions.
@@
First, to scale outbound communication from the exchange to a large number of participants, we employ an overlay multicast tree composed of a root exchange VM, proxy VMs as intermediate nodes, and participant VMs as leaf nodes. We develop a simple and effective heuristic to tune the tree’s fan-out and depth, navigating a tradeoff between increased serialization delay due to greater fan-out and increased propagation delay because of greater depth. To lower latency and counter the cloud’s variability, we pervasively employ hedging: routing redundant copies of the market data through multiple proxy and receiver VMs and rotating parent-to-child associations at each tree level on each multicasted packet.
Second, to scale inbound communication from participants to the exchange, we use the tree in reverse; this reduces the fan-in at the matching engine, decreasing incast-style drops. To further reduce drops during bursts of orders, every tree node paces inbound transmissions when the one-way delay between that node and its parent is high. While pacing reduces drops, it builds up queues. To counter this queuing latency, we develop a specialized scheduling algorithm, limit-order queuing (LOQ), which runs on the proxy VMs, and transmits market orders closest to the midpoint of the bid-ask spread—rather than transmitting FIFO. By prioritizing orders close to the midpoint, LOQ improves the likelihood that transmitted orders will be matched immediately. This improves matching throughput and order-matching latency relative to FIFO.
Third, we explore a range of trust models to capture how participants’ trading algorithms interface with the leaf nodes on the multicast tree. These include running the leaf within a VM and the algorithm within a protected enclave in the same VM, running the two entities as 2 separate cloud VMs, and running the 2 entities as 2 separate functions in a process. We describe and measure the performance vs. security tradeoffs inherent in each solution.
@@
Jasper can support a maximum throughput of 175K multicast messages per second at which point it is limited by a proxy VM’s egress bandwidth. It also scales to 1000 receivers, achieving a median multicast latency of while achieving a spatial variance of across these receivers. Jasper also scales better and achieves 50% lower latency compared to AWS’ Transit Gateway-based multicast. On the inbound side, Jasper efficiently handles large bursts of orders, using LOQ to nearly double the order-matching rate and maintain low order latencies during bursts.
To contextualize Jasper’s absolute performance, Jasper falls short of heavily engineered on-premises financial exchanges, which achieve spatial variance of a few nanoseconds [2]. Nevertheless, for certain financial instruments (e.g., cryptocurrency), their dominant exchanges [9, 10] have been built on the cloud since the beginning. For such instruments, Jasper is very competitive. Further, participant-to-exchange interfaces exist on a usability-performance tradeoff curve: colocated participants at one end and web APIs at the other [11, 12]. Jasper presents a compelling point on this curve: low and predictable latencies of ~ while scaling to 1K market participants and many additional benefits of the cloud.
2 Background
@@
The ME maintains a limit order book (LOB) (Fig. 3), which lists all bid and ask orders from MPs. When a bid price exceeds the lowest ask price, the orders are matched and executed. Unexecuted orders remain in the LOB, waiting for a match. In the price-time-priority algorithm [13], orders are arranged in price levels, with those at the same level sorted by arrival time, prioritizing older orders.111Price-time-priority is widespread, so we use it. Alternatives exist such as the pro-rata algorithm [13] An LOB snapshot shows a separation between bid and ask price levels, with the mid-price indicating the asset’s true value. Thus, orders closer to the mid-price have higher chances of getting matched early. The mid-price changes with new orders.
@@
Traditionally, the exchange and MP servers are colocated in the same datacenter and connected via equal-length wires using L1 [3] switches with switch multicast. The overall infrastructure is designed to provide low and equal latency from the exchange server to all MPs in both directions.
@@
Challenges of Cloud Migration. While the public cloud offers many advantages, it does not offer low-level control, e.g., allowing tenants to control wire lengths. The public cloud also exhibits high latency variance [14] and compute variance [15]. The latency between a pair of VMs can be significantly different from another VM pair [16]. Latency also fluctuates over time: Figure 1 plots the latency between a pair of VMs in an AWS region for a tumbling window of . The figure also shows infrequent but unpredictable latency spikes that increase latency substantially [14, 7]. Such phenomena make it challenging to achieve low and deterministic latency.
@@
Given the above challenges, it is reasonable to ask whether financial exchanges should ever be migrated to the cloud—and if so, how. This is an ongoing debate with reasonable arguments on both sides. Beyond the research projects [4, 7, 5] discussed below, some cloud providers are exploring close partnerships to build private clouds for exchanges’ bespoke requirements [17, 18]. Jasper contributes to this debate by empirically studying what performance guarantees can be achieved if a cloud tenant designs the exchange on the public cloud with publicly available cloud APIs. With Jasper’s DIY approach, a tenant does not need to wait for a cloud provider to build private clusters. While Jasper doesn’t achieve on-prem performance, its performance is competitive with several cloud-based exchanges [10, 9].
Prior Work and Jasper’s Motivation. CloudEx [4] synchronizes the clock of the exchange server VM to the clocks of several gateway VMs, to which participant VMs connect. On the outbound side, the exchange sets a release time for market data which the gateways respect to provide simultaneous delivery of market data. On the inbound side, each gateway timestamps its incoming participant’s orders; this timestamp determines an order’s priority in the price-time-priority algorithm. Octopus [5] implements an approach similar to CloudEx using exchange-controlled SmartNICs on the participant’s machines instead of gateway VMs to implement simultaneous release of market data. Recent literature including DBO [7], FBA [19], and Libra [20], proposes new definitions of fairness for exchanges.
Relative to these efforts, Jasper’s main contribution is architecting for scale. Jasper adopts the idea of synchronized clocks from CloudEx, but scales the system much further by using a communication tree in both the inbound and outbound directions and augmenting this tree with in-network pacing and scheduling algorithms (§3). Additionally, Jasper also tackles the latency-security tradeoff on the “last hop:” the link from the exchange’s edge (e.g., gateway VM, proxy, SmartNIC) to the participant’s trading algorithm (§6). Jasper employs the widely used notion of outbound fairness (simultaneous delivery of data to MPs) from CloudEx.
3 Jasper Overview
@@
Figure 3 presents an overview of Jasper. For communication between the exchange and participants, Jasper uses a bidirectional overlay tree: the exchange server is the root and MPs are the leaves, with intermediate proxy nodes. It is well known that trees help scale communication to many receivers [21, 22]. We build on such a tree to provide a multicast service for market data and handle bursty order submissions, but adapt the tree to the high-variance environment of the public cloud. Additionally, we adopt from CloudEx the idea of using synchronized clocks to (1) simultaneously disseminate market data and (2) sequence participants’ orders.
@@
Outbound: ME to Participants. On the outbound side of an exchange, we augment the base tree with 3 techniques to lower multicast latency and minimize the latency variance: (i) round-robin packet spraying (RRPS), (ii) proxy hedging and (iii) receiver hedging. All 3 hedge the risk of some part of the system showing performance variance, as explained below.
@@
To minimize the impact of latency spikes on the links between the tree nodes, proxies at each layer rotate their child VMs after each message, ensuring messages don’t always follow fixed paths. This approach limits the effect of latency spikes, as only a small subset of messages is impacted by any transient spike on any link within the tree.
@@
Inbound: Participants to ME. We use the same tree as above in the reverse direction for MP-to-exchange communication: MPs submit their orders to their parent proxies where it travels up the tree and reaches the ME at the root. Using a tree in the reverse direction reduces the fan-in factor of ME: the ME has to receive and process streams of orders from a small number of proxies instead of all MPs. This reduces packet losses as it reduces incast. We augment this reverse tree with in-network pacing and scheduling to efficiently handle bursty orders from MPs and relay them to the ME running at the root.
Packet pacing significantly reduces incast-style packet drops during periods of bursts. A proxy continuously senses One-Way Delay (OWD) between itself and its parent proxy. If OWD increases beyond a threshold, the proxy paces the packets to lower the OWD which in turn avoids packet drops.
However, pacing at proxies builds up queues, which are drained once a burst subsides. This can increase order latency and reduce ME throughput, which we compensate for via smarter scheduling. We develop Limit Order Queueing (LOQ) at each proxy, which prioritizes participant packets that contain critical orders, i.e., those closest to the mid-price. Critical orders are more likely to be matched upon reaching the ME, thus keeping the ME occupied without idling. Non-critical orders are likely to remain in the Limit Order Book (LOB) at the ME until the mid-price moves, and hence are deprioritized and wait longer at the proxies. LOQ increases the ME’s order matching rate and maintains low latency for matched orders during bursts.
Trust Models. Exchanges operate in an environment with strong incentives: participants have every incentive to act on market data as soon as it arrives and also have every incentive to set an earlier timestamp for their order. We develop multiple participant-to-exchange interfaces for the last hop between the participant’s trading algorithm and the leaf nodes of the tree. In the first interface, the leaf runs within a VM; the algorithm runs in a secure enclave within the VM. In the second, the leaf and algorithm run in separate VMs. In the third, the leaf and algorithm run in the same process.
These interfaces tradeoff performance vs. security: the first and second both don’t require the participant to reveal their algorithm to the exchange, but experience significant latency on the last hop. The third requires the participant to reveal their algorithm, at least in binary form, but has much lower last-hop latency. Between the first and second interfaces, we find that while their mean last hop latencies are comparable, the enclave-based solution has much lower variance.
4 Outbound Communication

### v11

@@
Abstract.
We present Jasper, a system for meeting the networking requirements of financial exchanges on the public cloud. Jasper uses an overlay tree to multicast market data from an exchange to 1000 participants with difference in data reception time between any two participants, crucial for maintaining fair competition. Jasper reuses the same tree for scalable inbound communication (participants to exchange), introducing a scheduling policy to enhance an exchange’s throughput during periods of bursty traffic. It also presents a message sequencing mechanism that achieves globally ordered delivery of messages from participants to the exchange, i.e., messages of participants are seen by the exchange in order they are generated to maintain fairness. Jasper achieves better scalability and 50% lower latency than the AWS multicast service [Services, ]. Jasper outperforms existing system, CloudEx [Ghalayini et al., 2021] in terms of supported number of participants, order matching rate of the exchange and multicast latency. Jasper’s techniques can be applied to other existing systems (e.g., DBO) to enhance their performance.
1. Introduction
@@
Financial exchanges are migrating to the public cloud for reasons such as improved scalability and reduced capital expenditure. Despite its benefits, the public cloud poses unique challenges. Exchanges have traditionally operated in on-premise or colocation facilities, engineered for deterministic and low latency. For fair market access, exchanges equalize cable lengths between the exchange and participant servers. This approach ensures that (i) all participants receive market data from the exchange simultaneously (outbound fairness) and (ii) an order generated earlier by one participant reaches the exchange before orders generated later by other participants (inbound fairness). It is common to employ low jitter switches (e.g., L1 switch [leptonsys.com, ]) to reduce latency variance.
@@
However, the public cloud lacks these enhancements. It is a best-effort environment characterized by nondeterminism (e.g. latency variance in Figure 1). In response, several projects have developed techniques for cloud-based exchanges. These include using synchronized clocks to compensate for nondeterminism [Ghalayini et al., 2021], using SmartNICs to hold data until all receiver has received it [Gong et al., 2024], and new fairness definitions [Goyal et al., 2022, Gupta et al., 2023]. These projects have demonstrated promising results for tens of participants but exhibit significant performance degradation as the number of participants increases. This limitation arises because scalability was not a primary design objective. Instead, the initial focus was on establishing a functional proof of concept for a fair exchange on the cloud. Having made significant progress in that regard, the next logical step is to consider scalability. Consequently, our paper develops techniques for scale: How do we architect an exchange to support communication between the exchange and ~1000 participants in the cloud while ensuring fairness in the exchange? In practice, a scalable exchange would also need scale-out compute techniques for the exchange server—a concern that is out of scope.
@@
Our system, Jasper, tackles two major challenges to provide network support for scalable exchanges. First, how do we support outbound communication of market data from the exchange to 1000 market participants, while ensuring (a) low spatial variance, i.e., all participants receive market data nearly simultaneously, (b) low latency from the exchange server to the participant, and (c) low temporal variance, i.e., latency doesn’t fluctuate over time? Second, how do we support inbound communication of participant orders to the exchange while (a) providing a fair chance of making trades to all the traders and, (b) achieving high throughput for the exchange server —especially during intense market activity when bursts of orders from many participants arrive at once, causing incast-style [Vasudevan et al., 2009] drops and increased latency? Jasper integrates multiple well-established mechanisms within a novel context as well as introduces new techniques to achieve high performance for a cloud-hosted exchange.
@@
First, to scale outbound communication from the exchange to a large number of participants, we employ an overlay multicast tree composed of a root exchange VM, proxy VMs as intermediate nodes, and participant VMs as leaf nodes. We develop a simple and effective heuristic to tune the tree’s fan-out and depth, navigating a trade-off between increased serialization delay due to greater fan-out and increased propagation delay because of greater depth. To lower latency and counter the cloud’s variability, we pervasively employ hedging: routing redundant copies of the market data through multiple proxy and receiver VMs and rotating parent-to-child associations at each tree level on each multicasted packet.
Second, to provide scalable and fair inbound communication from participants to the exchange, we propose a sequencer that relies on recent advancements in clock synchronization that is robust to latency fluctuations and works for VMs without hardware support. The sequencer ensures that an exchange server sees the messages generated by participants in their generation order so that inbound fairness is achieved regardless of the arbitrary message delays. We also propose a scheduling policy, Limit Order Queue, that helps achieve fairness and high order matching rate of the exchange during periods of bursty market activity. Finally, we reuse the overlay multicast tree in the reverse direction to proxy participants messages to the exchange which helps against incast-style packet drops by reducing the fan-in at the exchange when supporting a large number of participants. This leads to high throughput for the exchange server and low latency for trade orders.
@@
Jasper can support a maximum throughput of 175K multicast messages per second at which point it is limited by a proxy VM’s egress bandwidth. It also scales to 1000 receivers/VMs, achieving a median multicast latency of while maintaining a latency difference of across these receivers. Jasper also scales better and achieves 50% lower latency compared to AWS’ Transit Gateway-based multicast. On the inbound side, Jasper efficiently handles large bursts of orders, using LOQ to increase the order matching rate proportional to the bursts and maintain low order latencies. Jasper outperforms a prior system, CloudEx [Ghalayini et al., 2021] in terms of scalability, order matching rate and multicast latency. More importantly, the techniques described in Jasper are meant to provide a networking layer for the exchange systems and are thus composable with the existing exchange systems to enhance their performance. We think of Jasper as CloudEx augmented with several scaling methodologies that are able to maintain fairness. Onyx will be opensourced.
2. Background
@@
The ME maintains a limit order book (LOB) (Fig. 4), which lists all bid and ask orders from MPs. When a bid price exceeds or match the lowest ask price, the orders are executed (and matched together). Unexecuted orders remain in the LOB, waiting for a match. In the price-time-priority algorithm [investopedia.com, ],111Price-time-priority is widespread, so we use it. Alternatives exist such as the pro-rata algorithm [investopedia.com, ] orders are arranged in price levels, with those at the same level sorted by their arrival time at the exchange (which is equivalent to sorting them by their generation time at the MPs if all cables connecting MPs to exchange are precisely equalized). An LOB snapshot (Fig. 4) shows a separation between bid and ask price levels, with the mid-price indicating the asset’s true value. Thus, orders closer to the mid-price have higher chances of getting matched early. Any scheduling policy on orders should ensure that the semantics of price-time priority matching algorithm are not impacted i.e., the sequence of matched orders should remain unchanged.
@@
Traditionally, the exchange and MP servers are co-located in the same datacenter and connected via equal-length wires using L1 [leptonsys.com, ] switches with switch multicast. The overall infrastructure is designed to provide low and equal latency from the exchange server to all MPs in both directions.
@@
Challenges of Cloud Migration. While the public cloud offers many advantages, it does not offer low-level control, e.g., allowing tenants to control wire lengths. The public cloud also exhibits high latency variance [Hilyard et al., 2023] and computation variance [Venkataraman et al., 2016]. The latency between a pair of VMs can be significantly different from another pair of VMs [Vighnesh Sachidananda, ]. Latency also fluctuates over time: Figure 1 plots the latency between a pair of VMs in an AWS region for a tumbling window of . The figure also shows infrequent but unpredictable latency spikes that substantially increase latency [Hilyard et al., 2023, Gupta et al., 2023]. Such phenomena make it challenging to achieve low and deterministic latency in the public cloud.
@@
Given the above challenges, it is reasonable to ask whether financial exchanges should ever be migrated to the cloud — and if so, how. This is an ongoing debate with reasonable arguments on both sides. Beyond the research projects [Ghalayini et al., 2021, Gupta et al., 2023, Gong et al., 2024] discussed in the following, some cloud providers are exploring close partnerships to build private clouds for exchanges’ bespoke requirements [Group, , Osipovich, ]. Jasper contributes to this debate by studying what performance guarantees can be achieved if a cloud tenant designs the exchange architecture on the public cloud with publicly available cloud APIs. With Jasper’s DIY approach, a tenant does not need to wait for a cloud provider to build private clusters and can quickly ramp up a performant and fair exchange.
Is cloud migration still relevant? There has also been a debate on cloud migration and cost cuts being a paradox, giving emergence to cloud repatriation (moving away from cloud) as discussed by Martin Casado [a16z, ]. Cloud repatriation, unlike cloud migration, is highly relevant to large SaaS companies like Snowflake or Dropbox, which operate at a scale where they can benefit from economies of scale by owning infrastructure and have consistent workloads that justify such investments. However, this narrative doesn’t extend as well to smaller entities like financial institutions, which lack the consistent, cloud-scale workloads to make private infrastructure cost-effective and often benefit more from the flexibility and cost savings of not managing their own hardware.
Prior Work and Jasper’s Motivation. CloudEx [Ghalayini et al., 2021], the first cloud-hosted exchange system, may lose inbound as well as outbound fairness under latency fluctuations. To maintain fairness, participants wait until a set timeout (to counter latency fluctuations) before processing any order to ensure every participant has received the market data. The matching engine waits until a set timeout before processing any order to ensure that all the earlier generated orders have been received. As the number of participants increase, these timeouts need to increase (to not lose fairness) accordingly which lead to high latency and low order matching rate. DBO [Gupta et al., 2023] leverages mechanisms to always guarantee fairness among participants by decoupling fairness from latency fluctuations, although it is only applicable to a subset of trades that depend on the last received market data batch. As the number of participants increase, DBO also suffer from degraded performance because of (i) incast on the inbound side and, (ii) large latency of market data. Both of these systems can benefit from techniques to enhance the number of participants that maintain low latency, low latency fluctuations and high throughput for the exchange.
Consequently, Jasper’s main contribution is architecting for scale while maintaining fairness and achieving high performance. Jasper adopts the idea of synchronized clocks from CloudEx, but scales the system much further by using a communication tree in both the inbound and outbound directions to achieve high performance with a large number of participants. Jasper augments the tree with (i) a new sequencer to ensure inbound fairness under latency fluctuations (ii) a scheduling algorithm to achieve high performance under bursts of orders and, (iii) several performance variance mitigation techniques to help achieve fairness.
3. Jasper Overview
Setup and goals. Each MP VM hosts an order gateway and a trading algorithm.222“trading algorithm” and “MP” are interchangeably used Clocks of all MP VMs and the exchange server are synchronized using Huygens algorithm [Geng et al., 2018], which enables nanoseconds level synchronization for VMs without hardware support. Trading algorithm generates the orders and submit it to the colocated gateway. A gateway then attaches its current timestamp to an order and forwards it to the exchange (via TCP [Myers et al., 2024]), which hosts the matching engine (ME) to process orders. MP VMs are controlled by the exchange which loads the trading algorithms in them. Market data from the exchange (via UDP [Myers et al., 2024]) first arrives at a gateway which then forwards it to the trading algorithm. Such a model has been proposed previously [Haseeb et al., 2024, Yu et al., 2024].333An avid reader may be interested in a brief discussion of several deployment/trust models and their performance in Appendix G. Jasper aims to provide fairness during the competition among MPs for both the outbound (exchange to MPs) and inbound (MPs to exchange) directions. We specify the inbound/outbound fairness as follows.
Definition 3.1 (Outbound Fairness).
Every market data message sent from the exchange server to the market participants (MPs) should be seen by all the MPs simultaneously.
Definition 3.2 (Inbound Fairness).
An order generated earlier than other orders should be processed by the matching engine earlier than the other orders, irrespective of which MP generated which order.
@@ 3. Jasper Overview
While respecting the above fairness definitions, Jasper aims to achieve high performance i.e., supporting large number of MPs and achieving a high order matching rate, setting it apart, but complementary, from the previous work. Figure 2 presents an overview of Jasper architecture. Jasper employs a bidirectional overlay tree: the exchange server is the root and MPs are the leaves, with intermediate proxy nodes. It is well known that trees help scale communication to many receivers [Mokhtarian and Jacobsen, 2015, Rouskas and Baldine, 1997]. We build on such a tree to provide a multicast service for market data and handle order submissions from a large number of MPs, but adapt the tree to the high-variance environment of the public cloud.
@@
Outbound: ME to Participants (market data multicast). On the outbound side of an exchange, we augment the base tree with 3 techniques to lower multicast latency and minimize the latency variance: (i) round-robin packet spraying, (ii) proxy hedging and (iii) receiver hedging. All 3 hedge the risk of some part of the system exhibiting performance variance.
@@
To minimize the impact of latency spikes on the links between the tree nodes, proxies at each layer rotate their child VMs after each message, ensuring messages do not always follow fixed paths. This approach limits the effect of latency spikes, as only a small subset of messages is impacted by any transient spike on any link within the tree.
@@
Inbound: Participants to ME (orders submission). We develop a sequencer that sits at the ingress of the exchange server. All the incoming orders from MPs are fed to the sequencer while the output (the sequenced orders) are processed by the matching engine (ME) hosted by the exchange server. The sequencer ensures that inbound fairness is achieved regardless of latency fluctuations in the paths of orders. During periods of bursty market activity, MPs may generate a lot of orders overwhelming the ingress of the exchange. As we employ TCP in the inbound communication, the overwhelming of the exchange leads to queue build ups at the MP VMs. A special priority queue, Limit Order Queue (LOQ), runs at the egress of each MP VM for servicing the built up queues so that latency of orders remains low and high order matching rate at ME is achieved.
LOQ distinguishes critical orders –orders that will get matched by the ME upon their arrival at the exchange server– from non-critical orders– orders that will just stay in the limit order book and will not get immediately matched. LOQ prioritizes critical orders over noncritical orders so that ME processes the matchable orders earlier than non-matchable ones, leading to high order matching rate and low latency for orders. As we explain later, LOQ scheduling is designed to not affect inbound fairness.
Further, as the number of MPs increases, the ingress of the exchange server becomes a prominent bottleneck because of the large number of MPs simultaneously submitting orders. Even if the cumulative load of all the MPs is well under the ingress capacity of the exchange, the pattern of large number of orders from a large number of MPs simultaneously arriving at the exchange leads to incast-style packet drops. To resolve this, we reuse the proxy tree but in the reverse direction (for MP-to-exchange communication): MPs submit their orders to their parent proxies where it travels up the tree and reaches the ME at the root. Using a tree in the reverse direction reduces the fan-in factor of ME: the ME has to receive and process streams of orders from a small number of proxies instead of all MPs. This reduces packet losses, leads to fewer retransmissions by TCP which sums up to high throughput for the exchange server.
With the tree in place, we see queue build ups at the proxy nodes during bursty order arrivals. We apply our above-mentioned priority queue, LOQ, at the proxy nodes as well where it services the queues to significantly improve the order matching rate of the ME.
4. Market Data Multicast